OpenStudy (anonymous):

how can we sketch the graph r=3+sin theta

OpenStudy (amistre64):

ideally you pick a few points that are easy to identify, then plot them ....

OpenStudy (amistre64):

sin0=0, +3 sin45=sqrt2/2, +3 sin90 = 1, +3 etc ....